solve the equation 4s^2 + 36s + 81 = 0
Answer by nerdybill(7384)

solve the equation 4s^2 + 36s + 81 = 0
.
4s^2 + 36s + 81 = 0
4s^2 + 18s + 18s + 81 = 0
(4s^2 + 18s) + (18s + 81) = 0
2s(2s + 9) + 9(2s + 9) = 0
(2s + 9)(2s + 9) = 0
(2s + 9)^2 = 0
s = -9/2 = -4.5